# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ 6 $\overline{B}{}_{s}^{0}\to T^{0}_{c}\kappa^{0}$ $\frac{1}{\sqrt{2}}(C_{T}-C_{S}+C_{ST,1}-2C_{TS,2})$
$5$ $\overline{B}{}^{0}\to T^{+}_{c}\kappa^{-}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{ST,1}+C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2})$
$5$ $\overline{B}{}_{s}^{0}\to T^{+}_{c}a^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{S}+C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}+2E^{\prime}_{\bar{B}S,1}+E^{\prime}_{\bar{B}T})$
$6$ $B^{-}\to T^{0}_{c}\kappa^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{S}-C^{\prime}_{T}+C^{\prime}_{ST,2}-C^{\prime}_{TS,1})$
$6$ $\overline{B}{}^{0}\to T^{0}_{c}\overline{\kappa}{}^{0}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{S}-C^{\prime}_{T}-C^{\prime}_{ST,1}+2C^{\prime}_{TS,2})$
$6$ $\overline{B}{}_{s}^{0}\to T^{0}_{c}a_{0}^{0}$ $-\frac{1}{2}(C^{\prime}_{S}+C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}+2E^{\prime}_{\bar{B}S,1}+E^{\prime}_{\bar{B}T})$
$6$ $\overline{B}{}_{s}^{0}\to T^{0}_{c}f_{0}$ $\frac{c_{\phi}}{2}(C^{\prime}_{S}+2C^{\prime}_{T}-2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}+2E^{\prime}_{\bar{B}S,1}+E^{\prime}_{\bar{B}T})$
$-\frac{s_{\phi}}{\sqrt{2}}(C^{\prime}_{TS,1}+2E^{\prime}_{\bar{B}S,1}+E^{\prime}_{\bar{B}T})$
$6$ $\overline{B}{}_{s}^{0}\to T^{0}_{c}\sigma$ $-\frac{s_{\phi}}{2}(C^{\prime}_{S}+2C^{\prime}_{T}-2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}+2E^{\prime}_{\bar{B}S,1}+E^{\prime}_{\bar{B}T})$
$-\frac{c_{\phi}}{\sqrt{2}}(C^{\prime}_{TS,1}+2E^{\prime}_{\bar{B}S,1}+E^{\prime}_{\bar{B}T})$
$7$ $\overline{B}{}^{0}\to T^{\prime\prime+}_{c}\kappa^{-}$ $-\frac{1}{\sqrt{2}}(2T^{\prime}_{S}+C^{\prime}_{ST,1}+C^{\prime}_{ST,2}+C^{\prime}_{TS,1}+2C^{\prime}_{TS,2})$
$7$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime+}_{c}a^{-}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{S}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}+2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}-E^{\prime}_{\bar{B}T})$
$8$ $B^{-}\to T^{\prime\prime 0}_{c}\kappa^{-}$ $-\frac{1}{\sqrt{2}}(2T^{\prime}_{S}+C^{\prime}_{S}+C^{\prime}_{T}+C^{\prime}_{ST,2}+2C^{\prime}_{ST,3}+C^{\prime}_{TS,1})$
$8$ $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\overline{\kappa}{}^{0}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{S}+C^{\prime}_{T}-C^{\prime}_{ST,1}+2C^{\prime}_{ST,3}-2C^{\prime}_{TS,2})$
$8$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}a_{0}^{0}$ $-\frac{1}{2}(C^{\prime}_{S}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}+2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}-E^{\prime}_{\bar{B}T})$
$8$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}f_{0}$ $\frac{c_{\phi}}{2}(2C^{\prime}_{T}+C^{\prime}_{S}-2C^{\prime}_{ST,1}+C^{\prime}_{ST,2}+4C^{\prime}_{ST,3}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}$
$+2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}+3E^{\prime}_{\bar{B}T})+\frac{s_{\phi}}{\sqrt{2}}(-C^{\prime}_{TS,1}+2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}+E^{\prime}_{\bar{B}T})$
$8$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\sigma$ $-\frac{s_{\phi}}{2}(2C^{\prime}_{T}+C^{\prime}_{S}-2C^{\prime}_{ST,1}+C^{\prime}_{ST,2}+4C^{\prime}_{ST,3}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}$
$+2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}+3E^{\prime}_{\bar{B}T})+\frac{c_{\phi}}{\sqrt{2}}(-C^{\prime}_{TS,1}+2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}+E^{\prime}_{\bar{B}T})$
$9$ $\overline{B}{}_{s}^{0}\to T^{\prime\prime+}_{c\bar{s}}\kappa^{-}$ $-\frac{1}{\sqrt{2}}(2T^{\prime}_{S}+C^{\prime}_{S}+C^{\prime}_{ST,1}+2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}-E^{\prime}_{\bar{B}T})$
@@
The $\overline{B}\to TS$ decay amplitudes in scenario I, decomposed in topological amplitudes, are shown in Tables 15 and 16 for $\Delta S=0$ and $-1$ transitions, respectively, while those in scenario II are in Tables 17 and 18.
@@
In most of these cases, we need both $\Delta S=0$ and $-1$ transitions. Note that the last relations are interesting, because all three flavor exotic states are involved.
@@
$3^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}a_{0}^{0}$ $-\frac{1}{\sqrt{2}}(C_{ST,1}-C_{ST,2}-C_{TS,1})$
$3^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}f_{0}$ $-\frac{c_{\phi^{\prime}}}{\sqrt{2}}(2C_{T}+C_{ST,1}-C_{ST,2}+C_{TS,1})+s_{\phi^{\prime}}(C_{S}-C_{ST,1}-C_{TS,1})$
$3^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\sigma$ $\frac{s_{\phi^{\prime}}}{\sqrt{2}}(2C_{T}+C_{ST,1}-C_{ST,2}+C_{TS,1})+c_{\phi^{\prime}}(C_{S}-C_{ST,1}-C_{TS,1})$
$5^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{cs}\kappa^{0}$ $-C_{S}+C_{ST,1}+E_{\bar{B}T}$
$6^{\prime*}$ $B^{-}\to T^{-}_{cs}\kappa^{0}$ $-\sqrt{2}(C_{S}-C_{ST,1}+C_{ST,2})$
@@
$8^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{+}_{c\bar{s}\bar{s}}\kappa^{-}$ $-\sqrt{2}(C_{S}-C_{TS,1})$
@@
$11^{\prime*}$ $B^{-}\to T^{0}_{c}a_{0}^{-}$ $-\sqrt{\frac{2}{3}}(C_{T}-C_{ST,1}+C_{TS,1})$
@@
$11^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{c}\sigma$ $-\frac{s_{\phi}}{\sqrt{3}}(C_{T}+C_{ST,1}+C_{TS,1})+\sqrt{\frac{2}{3}}c_{\phi}(C_{T}-C_{S}+C_{ST,1}-C_{ST,2}+C_{TS,1})$
$11^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c}\kappa^{0}$ $-\sqrt{\frac{2}{3}}(C_{T}-C_{ST,2})$
@@
$12^{\prime*}$ $B^{-}\to T^{-}_{c}f_{0}$ $-c_{\phi^{\prime}}(C_{T}+C_{ST,1}+C_{TS,1})-\sqrt{2}s_{\phi}(C_{T}-C_{S}+C_{ST,1}-C_{ST,2}+C_{TS,1})$
$12^{\prime*}$ $B^{-}\to T^{-}_{c}\sigma$ $s_{\phi^{\prime}}(C_{T}+C_{ST,1}+C_{TS,1})-\sqrt{2}c_{\phi}C_{T}-C_{S}+C_{ST,1}-C_{ST,2}+C_{TS,1})$
@@
$14^{\prime}$ $\overline{B}{}^{\prime 0}\to T^{0}_{c}a_{0}^{0}$ $-\frac{1}{2\sqrt{6}}(C_{T}+C_{ST,1}-3C_{ST,2}-C_{TS,1}-E_{\bar{B}T})$
$14^{\prime}$ $\overline{B}{}^{0}\to T^{\prime 0}_{c}f_{0}$ $-\frac{c_{\phi}}{2\sqrt{6}}(C_{T}+C_{ST,1}-3C_{ST,2}+C_{TS,1}-3E_{\bar{B}T})$ $-\frac{s_{\phi}}{2\sqrt{3}}(-C_{S}+C_{T}+C_{ST,1}+2C_{ST,2}+C_{TS,1}+3E_{\bar{B}T})$
$14^{\prime}$ $\overline{B}{}^{0}\to T^{\prime 0}_{c}\sigma$ $\frac{s_{\phi}}{2\sqrt{6}}(C_{T}+C_{ST,1}-3C_{ST,2}+C_{TS,1}-3E_{\bar{B}T})$ $-\frac{c_{\phi}}{2\sqrt{3}}(-C_{S}+C_{T}+C_{ST,1}+2C_{ST,2}+C_{TS,1}+3E_{\bar{B}T})$
@@
$15^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime+}_{c\bar{s}}a_{0}^{-}$ $-\frac{1}{2}(2C_{S}-C_{ST,1}+C_{ST,2}-C_{TS,1})$
@@
$17^{\prime}$ $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}a_{0}^{0}$ $-\frac{1}{2\sqrt{2}}(C_{T}+C_{ST,1}+C_{ST,2}+4C_{ST,3}-C_{TS,1}-4E_{\bar{B}S,2}-E_{\bar{B}T})$
$17^{\prime}$ $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}f_{0}$ $-\frac{c_{\phi}}{2\sqrt{2}}(C_{T}+C_{ST,1}+C_{ST,2}+4C_{ST,3}+C_{TS,1}+4E_{\bar{B}S,2}+5E_{\bar{B}T})$
$+\frac{s_{\phi}}{2}(C_{S}-C_{T}-C_{ST,1}-2C_{ST,2}-4C_{ST,3}-C_{TS,1}-4E_{\bar{B}S,2}-3E_{\bar{B}T})$
@@
$2^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{+}_{c\bar{s}}\kappa^{-}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{S}-C^{\prime}_{ST,1}-E^{\prime}_{\bar{B}T})$
$3^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\overline{\kappa}{}^{0}$ $-C^{\prime}_{S}+C^{\prime}_{ST,1}+E^{\prime}_{\bar{B}T}$
$4^{\prime*}$ $\overline{B}{}^{0}\to T^{+}_{cs}a^{-}$ $-\sqrt{2}(C^{\prime}_{S}-C^{\prime}_{TS,1})$
@@
$5^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{cs}a_{0}^{0}$ $-\frac{1}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{S}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1})$
$5^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{cs}f_{0}$ $-\frac{c_{\phi}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{S}+2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}+C^{\prime}_{TS,1})-s_{\phi}(C^{\prime}_{T}+C^{\prime}_{TS,1})$
@@
$6^{\prime*}$ $B^{-}\to T^{-}_{cs}f_{0}$ $-c_{\phi}(C^{\prime}_{T}-C^{\prime}_{S}+2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}+C^{\prime}_{TS,1})-\sqrt{2}s_{\phi}(C^{\prime}_{T}+C^{\prime}_{TS,1})$
@@
$10^{\prime*}$ $\overline{B}{}^{0}\to T^{+}_{c}\kappa^{-}$ $-\sqrt{\frac{2}{3}}(C^{\prime}_{S}-C^{\prime}_{ST,1}+C^{\prime}_{ST,2})$

### v3

@@ 6 $\overline{B}{}_{s}^{0}\to T^{0}_{c}\kappa^{0}$ $\frac{1}{\sqrt{2}}(C_{T}-C_{S}+C_{ST,1}-2C_{TS,2})$
55 $\overline{B}{}^{0}\to T^{+}_{c}\kappa^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{ST,1}+C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2})$
55 $\overline{B}{}_{s}^{0}\to T^{+}_{c}a^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{S}+C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}+2E^{\prime}_{\bar{B}S,1}+E^{\prime}_{\bar{B}T})$
66 $B^{-}\to T^{0}_{c}\kappa^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{S}-C^{\prime}_{T}+C^{\prime}_{ST,2}-C^{\prime}_{TS,1})$
66 $\overline{B}{}^{0}\to T^{0}_{c}\overline{\kappa}{}^{0}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{S}-C^{\prime}_{T}-C^{\prime}_{ST,1}+2C^{\prime}_{TS,2})$
66 $\overline{B}{}_{s}^{0}\to T^{0}_{c}a_{0}^{0}$ $\frac{1}{2}(C^{\prime}_{S}+C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}+2E^{\prime}_{\bar{B}S,1}+E^{\prime}_{\bar{B}T})$
66 $\overline{B}{}_{s}^{0}\to T^{0}_{c}f_{0}$ $\frac{c_{\phi}}{2}(C^{\prime}_{S}+2C^{\prime}_{T}-2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}+2E^{\prime}_{\bar{B}S,1}+E^{\prime}_{\bar{B}T})$
$\frac{s_{\phi}}{\sqrt{2}}(C^{\prime}_{TS,1}+2E^{\prime}_{\bar{B}S,1}+E^{\prime}_{\bar{B}T})$
66 $\overline{B}{}_{s}^{0}\to T^{0}_{c}\sigma$ $\frac{s_{\phi}}{2}(C^{\prime}_{S}+2C^{\prime}_{T}-2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}+2E^{\prime}_{\bar{B}S,1}+E^{\prime}_{\bar{B}T})$
$\frac{c_{\phi}}{\sqrt{2}}(C^{\prime}_{TS,1}+2E^{\prime}_{\bar{B}S,1}+E^{\prime}_{\bar{B}T})$
77 $\overline{B}{}^{0}\to T^{\prime\prime+}_{c}\kappa^{-}$ $\frac{1}{\sqrt{2}}(2T^{\prime}_{S}+C^{\prime}_{ST,1}+C^{\prime}_{ST,2}+C^{\prime}_{TS,1}+2C^{\prime}_{TS,2})$
77 $\overline{B}{}_{s}^{0}\to T^{\prime\prime+}_{c}a^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{S}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}+2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}-E^{\prime}_{\bar{B}T})$
88 $B^{-}\to T^{\prime\prime 0}_{c}\kappa^{-}$ $\frac{1}{\sqrt{2}}(2T^{\prime}_{S}+C^{\prime}_{S}+C^{\prime}_{T}+C^{\prime}_{ST,2}+2C^{\prime}_{ST,3}+C^{\prime}_{TS,1})$
88 $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}\overline{\kappa}{}^{0}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{S}+C^{\prime}_{T}-C^{\prime}_{ST,1}+2C^{\prime}_{ST,3}-2C^{\prime}_{TS,2})$
88 $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}a_{0}^{0}$ $\frac{1}{2}(C^{\prime}_{S}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}+2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}-E^{\prime}_{\bar{B}T})$
88 $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}f_{0}$ $\frac{c_{\phi}}{2}(2C^{\prime}_{T}+C^{\prime}_{S}-2C^{\prime}_{ST,1}+C^{\prime}_{ST,2}+4C^{\prime}_{ST,3}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}$
$2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}+3E^{\prime}_{\bar{B}T})+\frac{s_{\phi}}{\sqrt{2}}(-C^{\prime}_{TS,1}+2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}+E^{\prime}_{\bar{B}T})$
88 $\overline{B}{}_{s}^{0}\to T^{\prime\prime 0}_{c}\sigma$ $\frac{s_{\phi}}{2}(2C^{\prime}_{T}+C^{\prime}_{S}-2C^{\prime}_{ST,1}+C^{\prime}_{ST,2}+4C^{\prime}_{ST,3}-C^{\prime}_{TS,1}-2C^{\prime}_{TS,2}$
$2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}+3E^{\prime}_{\bar{B}T})+\frac{c_{\phi}}{\sqrt{2}}(-C^{\prime}_{TS,1}+2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}+E^{\prime}_{\bar{B}T})$
99 $\overline{B}{}_{s}^{0}\to T^{\prime\prime+}_{c\bar{s}}\kappa^{-}$ $\frac{1}{\sqrt{2}}(2T^{\prime}_{S}+C^{\prime}_{S}+C^{\prime}_{ST,1}+2E^{\prime}_{\bar{B}S,1}+2E^{\prime}_{\bar{B}S,2}-E^{\prime}_{\bar{B}T})$
@@
The $\overline{B}\to TS$ decay amplitudes in scenario I, decomposed in topological amplitudes, are shown in Tables 15 and 16 for $\Delta S=0$ and 1-1 transitions, respectively, while those in scenario II are in Tables 17 and 18.
@@
In most of these cases, we need both $\Delta S=0$ and 1-1 transitions. Note that the last relations are interesting, because all three flavor exotic states are involved.
@@
$3^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}a_{0}^{0}$ $\frac{1}{\sqrt{2}}(C_{ST,1}-C_{ST,2}-C_{TS,1})$
$3^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}f_{0}$ $\frac{c_{\theta}}{\sqrt{2}}(2C_{T}+C_{ST,1}-C_{ST,2}+C_{TS,1})+s_{\theta}(C_{S}-C_{ST,1}-C_{TS,1})$
$3^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\sigma$ $\frac{s_{\theta}}{\sqrt{2}}(2C_{T}+C_{ST,1}-C_{ST,2}+C_{TS,1})+c_{\theta}(C_{S}-C_{ST,1}-C_{TS,1})$
$5^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{cs}\kappa^{0}$ $C_{S}+C_{ST,1}+E_{\bar{B}T}$
$6^{\prime*}$ $B^{-}\to T^{-}_{cs}\kappa^{0}$ $\sqrt{2}(C_{S}-C_{ST,1}+C_{ST,2})$
@@
$8^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{+}_{c\bar{s}\bar{s}}\kappa^{-}$ $\sqrt{2}(C_{S}-C_{TS,1})$
@@
$11^{\prime*}$ $B^{-}\to T^{0}_{c}a_{0}^{-}$ $\sqrt{\frac{2}{3}}(C_{T}-C_{ST,1}+C_{TS,1})$
@@
$11^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{c}\sigma$ $\frac{s_{\phi}}{\sqrt{3}}(C_{T}+C_{ST,1}+C_{TS,1})+\sqrt{\frac{2}{3}}c_{\phi}(C_{T}-C_{S}+C_{ST,1}-C_{ST,2}+C_{TS,1})$
$11^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c}\kappa^{0}$ $\sqrt{\frac{2}{3}}(C_{T}-C_{ST,2})$
@@
$12^{\prime*}$ $B^{-}\to T^{-}_{c}f_{0}$ $c_{\theta}(C_{T}+C_{ST,1}+C_{TS,1})-\sqrt{2}s_{\phi}(C_{T}-C_{S}+C_{ST,1}-C_{ST,2}+C_{TS,1})$
$12^{\prime*}$ $B^{-}\to T^{-}_{c}\sigma$ $s_{\theta}(C_{T}+C_{ST,1}+C_{TS,1})-\sqrt{2}c_{\phi}C_{T}-C_{S}+C_{ST,1}-C_{ST,2}+C_{TS,1})$
@@
$14^{\prime}$ $\overline{B}{}^{\prime 0}\to T^{0}_{c}a_{0}^{0}$ $\frac{1}{2\sqrt{6}}(C_{T}+C_{ST,1}-3C_{ST,2}-C_{TS,1}-E_{\bar{B}T})$
$14^{\prime}$ $\overline{B}{}^{0}\to T^{\prime 0}_{c}f_{0}$ $\frac{c_{\phi}}{2\sqrt{6}}(C_{T}+C_{ST,1}-3C_{ST,2}+C_{TS,1}-3E_{\bar{B}T})$ $\frac{s_{\phi}}{2\sqrt{3}}(-C_{S}+C_{T}+C_{ST,1}+2C_{ST,2}+C_{TS,1}+3E_{\bar{B}T})$
$14^{\prime}$ $\overline{B}{}^{0}\to T^{\prime 0}_{c}\sigma$ $\frac{s_{\phi}}{2\sqrt{6}}(C_{T}+C_{ST,1}-3C_{ST,2}+C_{TS,1}-3E_{\bar{B}T})$ $\frac{c_{\phi}}{2\sqrt{3}}(-C_{S}+C_{T}+C_{ST,1}+2C_{ST,2}+C_{TS,1}+3E_{\bar{B}T})$
@@
$15^{\prime}$ $\overline{B}{}_{s}^{0}\to T^{\prime+}_{c\bar{s}}a_{0}^{-}$ $\frac{1}{2}(2C_{S}-C_{ST,1}+C_{ST,2}-C_{TS,1})$
@@
$17^{\prime}$ $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}a_{0}^{0}$ $\frac{1}{2\sqrt{2}}(C_{T}+C_{ST,1}+C_{ST,2}+4C_{ST,3}-C_{TS,1}-4E_{\bar{B}S,2}-E_{\bar{B}T})$
$17^{\prime}$ $\overline{B}{}^{0}\to T^{\prime\prime 0}_{c}f_{0}$ $\frac{c_{\phi}}{2\sqrt{2}}(C_{T}+C_{ST,1}+C_{ST,2}+4C_{ST,3}+C_{TS,1}+4E_{\bar{B}S,2}+5E_{\bar{B}T})$
$\frac{s_{\phi}}{2}(C_{S}-C_{T}-C_{ST,1}-2C_{ST,2}-4C_{ST,3}-C_{TS,1}-4E_{\bar{B}S,2}-3E_{\bar{B}T})$
@@
$2^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{+}_{c\bar{s}}\kappa^{-}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{S}-C^{\prime}_{ST,1}-E^{\prime}_{\bar{B}T})$
$3^{\prime*}$ $\overline{B}{}_{s}^{0}\to T^{0}_{c\bar{s}}\overline{\kappa}{}^{0}$ $C^{\prime}_{S}+C^{\prime}_{ST,1}+E^{\prime}_{\bar{B}T}$
$4^{\prime*}$ $\overline{B}{}^{0}\to T^{+}_{cs}a^{-}$ $\sqrt{2}(C^{\prime}_{S}-C^{\prime}_{TS,1})$
@@
$5^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{cs}a_{0}^{0}$ $\frac{1}{\sqrt{2}}(C^{\prime}_{T}+C^{\prime}_{S}-C^{\prime}_{ST,2}-C^{\prime}_{TS,1})$
$5^{\prime*}$ $\overline{B}{}^{0}\to T^{0}_{cs}f_{0}$ $\frac{c_{\phi}}{\sqrt{2}}(C^{\prime}_{T}-C^{\prime}_{S}+2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}+C^{\prime}_{TS,1})-s_{\phi}(C^{\prime}_{T}+C^{\prime}_{TS,1})$
@@
$6^{\prime*}$ $B^{-}\to T^{-}_{cs}f_{0}$ $c_{\phi}(C^{\prime}_{T}-C^{\prime}_{S}+2C^{\prime}_{ST,1}-C^{\prime}_{ST,2}+C^{\prime}_{TS,1})-\sqrt{2}s_{\phi}(C^{\prime}_{T}+C^{\prime}_{TS,1})$
@@
$10^{\prime*}$ $\overline{B}{}^{0}\to T^{+}_{c}\kappa^{-}$ $\sqrt{\frac{2}{3}}(C^{\prime}_{S}-C^{\prime}_{ST,1}+C^{\prime}_{ST,2})$